OpenStudy (jazzyfa30):

Use the substitution method to solve: m - n = 7 3m + 2n = 6. Select one: a. (-3,-10) b. (4.-3) c. (-5,2) d. (-4,3)

OpenStudy (jazzyfa30):

@phi help

OpenStudy (anonymous):

using first equation find m in terms of n m=7+n put this value of m in the second equation 3(7+n)+2n=6 simplify and find n and then use m=7+n to find m